# Acoustoelectric Amplification in a Piezoelectric-2DEG Heterostructure

### v1

@@ Abstract
We provide a quantum mechanical description of phonon amplification in a heterostructure consisting of a two-dimensional electron gas (2DEG) stacked on top of a piezoelectric material. An applied drift voltage effectively creates a population inversion in the momentum states of the 2DEG electrons, giving rise to spontaneous emission of phonons. Once an acoustic wave is launched, the pumped electrons release phonons via stimulated emission, returning to depleted ground states before being pumped back to the excited states. We show that whereas efficient amplification using a 1D electron gas requires the acoustic wavelength to roughly equal the average electron-electron spacing, a 2DEG enables efficient amplification for any wavelength greater than the average electron-electron spacing. We derive the imaginary and real parts of the 2DEG’s first-order acoustic susceptibility as functions of electronic drift velocity in specific limits and derive the gain per unit length for the signal and the quantum noise, with the gain matching the classical result in the short-electronic-lifetime (low-mobility) regime. Moreover, we analyze the gain clamping due to pump depletion and calculate the maximum achievable intensity. Our results provide a framework for designing novel acoustic devices including a quantum phononic laser and phase-insensitive quantum phononic amplifiers.
@@ III.5 High-Drift-Velocity Regime
The real and imaginary parts of the susceptibility in the high-drift-velocity limit are shown in Fig. 7. Note that these are in log-log scale, with the first plot actually depicting the opposite of the real susceptibility (since the real susceptibility in this regime is negative). As expected, both the real and imaginary parts decline with increasing drift velocity in this regime. However, while the real part for any given drift velocity is invariant in mobility, the imaginary part is about 10 times lower for a mobility of $1000$ compared to $100\textrm{ m}^{2}/(\textrm{V}\cdot\textrm{s})$, reflecting the aforementioned inverse variation between the imaginary susceptibility and the mobility.
@@ V Clamping
Having determined the amplifier gain, we now shift our focus to deriving the clamping condition. Here, it is essential to account for the general time-evolution of the phonon mode through a beam-splitter Hamiltonian covering both the phonon and the 2DEG electron modes. Assuming a beam-splitter Hamiltonian consisting of $N$ pumped electrons each coupling with the phonon field through a time-varying coupling coefficient $g_{j}(t)$ (corresponding to the $j^{\textrm{th}}$ electron), the noise and the signal intensity feature the same gain per unit length (as shown in Appendix D), causing the signal-to-noise ratio to remain constant throughout the amplification process (analogously to a laser or maser). This common gain per unit length is given in terms of the coupling coefficient between a single electron and a phonon as follows:
@@
where $\tau$ is the time elapsed in a single amplifier segment (with a corresponding length of $v_{s}\tau$), $g=\left\langle\Big|\left\langle g_{j}(t)\right\rangle_{t}\Big|^{2}\right\rangle_{j}^{1/2}$ is root-mean-squared (over the $N$ electrons) of the time-averaged amplitude (averaged over the period $\tau$) of the coupling coefficient, and $M=L/(v_{s}\tau)$ is the number of amplifier segments. We can conceptualize $\tau$ as the effective reset time for an electron after it emits a phonon (i.e., the time the electron takes to return to the original state). Given an electronic decay rate $\gamma/2$ for the initial and final states each, the overall reset time is given approximately as $\tau\approx 1/\gamma$.
Although the gain per unit length at low phonon intensity is proportional to $g^{2}N$, the clamping condition will be solely dependent on $g$ itself, as we will show later in this section. As such, we aim to dis-aggregate $g$ and $N$ by separately deriving $g$. We start by solving the interaction coefficient $g_{\mathrm{emit}}$ corresponding to phonon emission by pumped electrons. This coefficient acquires a time-dependence due to electronic decay and electron-phonon detuning, as shown in the following composite interaction Hamiltonian, where we sum over the interactions with all electrons in the region $S$ (see Fig. 3(a)):
@@ Acknowledgements.
This work was supported by the Laboratory Directed Research and Development program at Sandia National Laboratories, a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia LLC, a wholly owned subsidiary of Honeywell International Inc. for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525. This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. DOE’s National Nuclear Security Administration under contract DE-NA-0003525. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government.
@@ Appendix D Amplified Phonon Wave and Quantum Noise
Note that all $N$ electrons can potentially change to $|l\rangle_{k}$ state while emitting a phonon with an energy $\hbar\omega$. To be clear, we will divide the entire amplifier length into $M$ segments with identical lengths. $N$ is the number of “excited” electrons in each segment.
A quantum operator $A$ evolves in time as
@@
We start with the following beam-splitter Hamiltonian, consisting of a sum over $N$ electron-phonon interactions, each with a time-varying coupling coefficient $g_{j}(t)$:
@@
where $G_{0}=1+(g\tau)^{2}N$, and $g$ is defined as the square-root of the expectation value of $|g_{j}|^{2}$ over the $N$ electrons, i.e., $g=\left\langle|g_{j}|^{2}\right\rangle_{j}^{1/2}=\left\langle\Big|\left\langle g_{j}(t)\right\rangle_{t}\Big|^{2}\right\rangle_{j}^{1/2}$. We also take the partial expected operator of the phonon field as:

### v2

# Quantum-Limited Acoustoelectric Amplification in a Piezoelectric-2DEG Heterostructure

@@ Abstract
We provide a quantum mechanical description of phonon amplification in a heterostructure consisting of a two-dimensional electron gas (2DEG) stacked on top of a piezoelectric material. An applied drift voltage effectively creates a population inversion in the momentum states of the 2DEG electrons, giving rise to spontaneous emission of phonons. Once an acoustic wave is launched, the pumped electrons release phonons via stimulated emission, returning to depleted ground states before being pumped back to the excited states. We show that whereas efficient amplification using a 1D electron gas requires the acoustic wavelength to roughly equal the mean distance between electrons, a 2DEG enables efficient amplification for any wavelength greater than this distance. We derive the imaginary and real parts of the 2DEG’s first-order acoustic susceptibility as functions of electronic drift velocity in specific limits and derive the gain per unit length for the signal and the quantum noise, with the gain matching the classical result in the short-electronic-lifetime (low-mobility) regime. Moreover, we analyze the gain clamping due to pump depletion and calculate the maximum achievable intensity. Our results provide a framework for designing novel acoustic devices including a quantum phononic laser and phase-insensitive quantum phononic amplifiers.
@@ III.5 High-Drift-Velocity Regime
The real and imaginary parts of the susceptibility in the high-drift-velocity limit are shown in Fig. 7. Note that these are in log-log scale, with the first plot actually depicting the opposite of the real susceptibility (since the real susceptibility in this regime is negative). As expected, both the real and imaginary parts decline with increasing drift velocity in this regime. However, while the real part for any given drift velocity is invariant in mobility, the imaginary part is about 10 times lower for a mobility of 10001000 compared to $100\textrm{ m}^{2}/(\textrm{V}\cdot\textrm{s})$, reflecting the aforementioned inverse variation between the imaginary susceptibility and the mobility.
@@ V Clamping
Having determined the amplifier gain, we now shift our focus to deriving the clamping condition. Here, it is essential to account for the general time-evolution of the phonon mode through a beam-splitter Hamiltonian covering both the phonon and the 2DEG electron modes. Assuming a beam-splitter Hamiltonian consisting of NN pumped electrons each coupling with the phonon field through a time-varying coupling coefficient $g_{j}(t)$ (corresponding to the $j^{\textrm{th}}$ electron), the noise and the signal intensity feature the same gain per unit length (as shown in Appendix D), causing the signal-to-noise ratio to remain constant throughout the amplification process (analogously to a laser or maser). This common gain per unit length is given in terms of the coupling coefficient between a single electron and a phonon as follows:
@@
where $\tau$ is the time elapsed in a single amplifier segment (with a corresponding length of $v_{s}\tau$), $g=\left\langle\Big|\left\langle g_{j}(t)\right\rangle_{t}\Big|^{2}\right\rangle_{j}^{1/2}$ is root-mean-squared (over the NN electrons) of the time-averaged amplitude (averaged over the period $\tau$) of the coupling coefficient, and $M=L/(v_{s}\tau)$ is the number of amplifier segments. We can conceptualize $\tau$ as the effective reset time for an electron after it emits a phonon (i.e., the time the electron takes to return to the original state). Given an electronic decay rate $\gamma/2$ for the initial and final states each, the overall reset time is given approximately as $\tau\approx 1/\gamma$.
Although the gain per unit length at low phonon intensity is proportional to $g^{2}N$, the clamping condition will be solely dependent on $g$ itself, as we will show later in this section. As such, we aim to dis-aggregate $g$ and NN by separately deriving $g$. We start by solving the interaction coefficient $g_{\mathrm{emit}}$ corresponding to phonon emission by pumped electrons. This coefficient acquires a time-dependence due to electronic decay and electron-phonon detuning, as shown in the following composite interaction Hamiltonian, where we sum over the interactions with all electrons in the region $S$ (see Fig. 3(a)):
@@ Acknowledgements.
This work was supported by the Laboratory Directed Research and Development program at Sandia National Laboratories. This work was also performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. DOE’s National Nuclear Security Administration under contract DE-NA-0003525. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government. In addition, this material is based on research sponsored in part by the Defense Advanced Research Projects Agency (DARPA) through a Young Faculty Award (YFA) under grant D23AP00174-00 and through DARPA contract DARPA-PA-23-03-01. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, by DARPA, the Department of the Interior, or the US Government.
@@ Appendix D Amplified Phonon Wave and Quantum Noise
Note that all NN electrons can potentially change to $|l\rangle_{k}$ state while emitting a phonon with an energy $\hbar\omega$. To be clear, we will divide the entire amplifier length into $M$ segments with identical lengths. NN is the number of “excited” electrons in each segment.
A quantum operator AA evolves in time as
@@
We start with the following beam-splitter Hamiltonian, consisting of a sum over NN electron-phonon interactions, each with a time-varying coupling coefficient $g_{j}(t)$:
@@
where $G_{0}=1+(g\tau)^{2}N$, and $g$ is defined as the square-root of the expectation value of $|g_{j}|^{2}$ over the NN electrons, i.e., $g=\left\langle|g_{j}|^{2}\right\rangle_{j}^{1/2}=\left\langle\Big|\left\langle g_{j}(t)\right\rangle_{t}\Big|^{2}\right\rangle_{j}^{1/2}$. We also take the partial expected operator of the phonon field as: